OpenStudy (anonymous):

Before beginning a fundraiser, Lindsey estimated that the fundraiser would earn $500. She sold 114 raffle tickets at $5 each for the fundraiser, actually earning $5 • 114 = $570 for the fundraiser. What was the percent error in Lindsey’s estimate? Round to the nearest tenth of a percent, if necessary. A. 12.3% B. 14% C. 70% D. 77.2%

OpenStudy (anonymous):

Help please, I'll medal and fan!!

OpenStudy (anonymous):

The answer would be B. 14% Take 570 - 500 = 70 Do 70/500 * (100) to get percentage So we get 0.14 * 100 = 14% which is B.

OpenStudy (anonymous):

Percent error is the difference between the actual value and the real value divided by the real value and multiply by 100. So to make this easier: \[\frac{ \left| actual-(measured/guessed/calculated) \right| }{ real value }\]

OpenStudy (anonymous):

He's wrong its 12.3%

OpenStudy (anonymous):

amberosales is correct it is 12.3% for new comers to this question